OpenStudy (anonymous):

Solve the logarithmic equation: _4 means log base 4 (for clarification) (a) log_4 (2x-18) - log_4 (x-25) = 1 (b) log_4 (x-18) + log_4 (x-3) = 2

geerky42 (geerky42):

ln(a) + ln(b) = ln(ab) ln(a) - ln(b) = ln(a/b)

OpenStudy (anonymous):

Hmm... How did you change it to ln instead of log?

geerky42 (geerky42):

You don't have to change it. ln is log_e. It works with all logarithm; bases have to be same, though.

geerky42 (geerky42):

In your problem, bases are same, so use these properties.

OpenStudy (anonymous):

So log_4(2x-18)(x-25)? wouldn't that factor out all crazy and not equal 1?

OpenStudy (anonymous):

oh I got messed up. that would be... log_4 (2x-18)/log_4 (x-25) = 1 Then what I drop the log_4 since they cancel out?

geerky42 (geerky42):

log_4 (2x-18)/(x-25) = 1 It looks ugly but it will look nicer when you convert this equation into exponent form.

geerky42 (geerky42):

Can you do it?

OpenStudy (anonymous):

no you lost me. I was absent that day and didn't pick up any of this. I know it's from Algebra but that has been a very long time ago.

geerky42 (geerky42):

BASE^EXPONENT = ANSWER log_BASE ANSWER = EXPONENT

geerky42 (geerky42):

For explain, \(\large log_5 7x^2 = 4 \Rightarrow 5^4 = 7x^2\). Still lost?

OpenStudy (anonymous):

Yea lost. I am trying to look up other references cause I am completely confused.

geerky42 (geerky42):

This may help you: http://www.youtube.com/watch?v=yLaJydYp4BQ

jimthompson5910 (jim_thompson5910):

log_4 (2x-18) - log_4 (x-25) = 1 log_4 [ (2x-18)/(x-25) ] = 1 (2x-18)/(x-25) = 4^1 (2x-18)/(x-25) = 4 2x-18 = 4(x-25) I'll let you take it from here. Remember to check any and all possible answers

OpenStudy (anonymous):

still lost.

geerky42 (geerky42):

Did you watch the video from the given link above here?

OpenStudy (anonymous):

yea

geerky42 (geerky42):

Still lost? Try another video on suggestions list.

geerky42 (geerky42):